(1 point) At a certain university, the Math SAT score and first-year GPA for the students that just completed their freshman year have the following summary statistics (if you cannot see the table, go to User Settings and Change Display Options to images): Average Math SAT score ะ 540 SD 100 A GPA 2.8 SD 0.6 verage first-vear r 0.58 The scatterplot of first-year GPAs on Math SAT scores is an oval-shaped cloud of points. Please give your answers correct to four decimal places. 1. Predict the first-year GPA of one of these students that got a 700 on the Math SAT. 2. Predict the first-year GPA of one of these students that got a 400 on the Math SAT 3. Predict the Math SAT score of a student that has a first-year GPA of 3.8. 4. Predict the Math SAT score of a student that has a first-year GPA of 2.0. 5. Fill in the blank: According to the regression method, every 10 points on the Math SAT is worth about points in a students GPA, on average.
Answer #1

1) predicted GPA =2.8+0.58*(700-540)*0.6/100=3.3568

2)predicted GPA =2.8+0.58*(400-540)*0.6/100=2.3128

3)predicted Math SAT score =540+0.58*(3.8-2.8)*100/0.6=636.6667

4)

predicted Math SAT score =540+0.58*(2-2.8)*100/0.6=462.6667

5)

worth about (0.58*10*0.6/100)=0.00348 points in a students GPA